A wire of length 1 m and radius 0.1 mm has a resistance of 100 Ω. Find the resistivity of the material.

Grade:upto college level

1 Answers

Aditi Chauhan
9 years ago
Sol. ℓ = 1 m, r = 0.1 mm = 0.1 */ 10^–3 m R = 100 Ω, f = ? ⇒ R = f ℓ / a ⇒ f = Ra/ℓ = 100 * 3.14 * 0.1 * 0.1 *10^-6/1 = 3.14 * 10^-6 = π * 10^-6 Ω-m.